OpenStudy (anonymous):

Write the quadratic function in vertex form. y = x2 + 4x + 7

OpenStudy (anonymous):

this is pre cal

OpenStudy (anonymous):

can you complete the square?

OpenStudy (jhannybean):

Yes.

OpenStudy (anonymous):

i forgot how to

OpenStudy (anonymous):

alright... x^2 +4x_____)+7 divide 4 by 2 and square it = 4 now factor (x^2 +4x +4)+7

OpenStudy (jhannybean):

You've got \[\large y=x^2 +4x+7\]\[\large y=(x^2+4x) +7\]\[\large y= (x^2 +4x+4 -4) +7\]\[\large y= (x+2)^2 +3\]

OpenStudy (anonymous):

I mean x^2 +4x +4) +7 -4

OpenStudy (jhannybean):

Vertex form follows the format \(\large y=a(x-h)^2 +k\) so just match up \[\large y=\color{green}1(x-\color{blue}{(-2)})^2 +\color{blue}3\] and you'll find your vertex and all that other good stuff.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

tnks

OpenStudy (jhannybean):

no problemo.